### v2

# Model Inversion Attacks Meet Cryptographic Fuzzy Extractors

Abstract
Model inversion attacks pose an open challenge to privacy-sensitive applications that use machine learning (ML) models. For example, face authentication systems use modern ML models to compute embedding vectors from face images of the enrolled users and store them. If leaked, inversion attacks can accurately reconstruct user faces from the leaked vectors. There is no systematic characterization of properties needed in an ideal defense against model inversion, even for the canonical example application of a face authentication system susceptible to data breaches, despite a decade of best-effort solutions.
In this paper, we formalize the desired properties of a provably strong defense against model inversion and connect it, for the first time, to the cryptographic concept of fuzzy extractors. We further show that existing fuzzy extractors are insecure for use in ML-based face authentication. We do so through a new model inversion attack called Pipe, which achieves a success rate of over $89\%$ in most cases against prior schemes. We then propose L2FE-Hash, the first candidate fuzzy extractor which supports standard Euclidean distance comparators as needed in many ML-based applications, including face authentication. We formally characterize its computational security guarantees, even in the extreme threat model of full breach of stored secrets, and empirically show its usable accuracy in face authentication for practical face distributions. It offers attack-agnostic security without requiring any re-training of the ML model it protects. Empirically, it nullifies both prior state-of-the-art inversion attacks as well as our new Pipe attack.
[…]
1 Introduction
Machine learning (ML) systems are increasingly being used in security-sensitive applications, such as for authenticating users [2022:arcface, 2015:facenet], establishing content ownership [2021:neural-hash], CSAM detection [2021:neural-hash], and so on. In these applications, neural networks often encode security-sensitive inputs into vector representations called embedding vectors. For example, in face authentication, standard neural networks generate embeddings corresponding to a user’s face image [2022:arcface, 2015:facenet]. These embedding vectors are stored on databases of application service providers [2023:clearview-ai] or on authentication devices [2024:biometric-wallet].
Databases and devices storing sensitive information are frequently prone to breaches [2024:outabox-breach] and theft/loss [2024:biometric-wallet] respectively. This gives rise to a threat model where the attacker gains access to all information persistently stored in the database or the device, as is typical of attacks colloquially referred to “smash-and-grab” attacks. The breached information can be used to recover the original security-sensitive data corresponding to the embedding vectors, causing severe privacy risks. A well-known class of such attacks is called model inversion. Since their discovery in 2015 on deep networks [2015:model-inversion], no defenses with strong formal security have yet been developed for model inversion attacks.
Model inversion attacks allow the adversary to recover inputs, given the outputs of the ML model. The canonical example introduced by the original work of Fredrikson et al. showcases the threat to face recognition systems, where outputs of deep networks can be used to recover the input face images reasonably accurately [2015:model-inversion]. Over the last decade, ML models have evolved to serve tasks beyond classification which motivated the original work of Fredrikson et al. The outputs of modern ML models are embedding vectors which can be compared in $\ell_{2}$ distance to other vectors. These vectors can be used to implement classification in a closed-set [2021:text-embedding-classification], similarity testing in an open-set [2018:LPIPS, 2015:facenet], lookup in RAG databases [2024:rag-databases], or generative tasks like text-to-image generation [2021:text-to-image]. Recent works have shown that embedding vectors generated from today’s ML models, if stored unprotected (in the clear), are highly susceptible to inversion attacks [2023:stylegan-recontruction-attack, 2023:id3pm-diffusion-inversion-attack, 2024:bob]. Therefore, it is important to devise a generic defense for model inversion attacks that works for applications where vector comparison with the standard Euclidean ($\ell_{2}$) distance is used. In this paper, we will use the classical example of face recognition ML models, but our goal is to motivate a broader defense strategy.
Cryptographic post-processing defenses.
One way to build a generic defense with provable guarantees is to turn to cryptographic methods. Specifically, can one post-process the output of the ML model with a cryptographic hash [^1] If we encrypt the embeddings instead, the encryption key must be persistently stored. This offers no security when all persistent data is leaked. before they are stored, to create what we call protected embeddings, such that they can still be used for $\ell_{2}$-based authentication while resisting model inversion attacks? The advantage of post-processing defenses is that they do not require re-training of the ML model, and they are agnostic to any specific model inversion attack. They offer a complementary approach to many in-training defense methods [2022:bido, 2021:mid, 2024:midre, 2023:defense-gan], which offer no formal security guarantees, nor do they aim to develop an attack-agnostic defense. Remarkably, despite their appeal, post-processing defenses have not yet been considered for tackling model inversion.
To understand why, it is perhaps useful to draw parallels with how hashing protects against the same threat model of password database breaches [2019:ml-crack-password]. It is standard practice to store passwords salted-and-hashed using standard cryptographic hash functions like SHA256, so that inversion of the plaintext passwords from a leaked password database becomes hard [1979:salt-and-hash]. However, such hash functions do not work for protecting embeddings generated from ML models. This is because ML models are designed to tolerate noise: Embeddings generated from two different face images are considered to be of the same person if their generated embedding vectors are close in $\ell_{2}$ distance. In face authentication, the user being authenticated must have an embedding vector that is close to the one expected, unlike passwords that must match exactly. However, standard cryptographic hash functions (eg. SHA256) used for salting-and-hashing passwords cannot tolerate any errors in inputs, i.e., even a single bit change in input induces a large change in output of the hash function. Hence, standard hash functions like SHA256 cannot be used to protect ML embedding vectors.
Figure 1: An example inverting the FacialFE-protected embedding vector to obtain a face image similar to the original using Pipe.
Post-processing defenses for tackling model inversion attacks on modern ML models have not been demonstrated so far. Nonetheless, there have been partially-secure post-processing candidates that one could leverage from prior literature as a starting point [2011:survey-biocrypto, 2018:survey-biocrypto]. These schemes allow noisy inputs to authenticate successfully as desired. Indeed, in our evaluation, we show that these schemes are effective against prior model inversion attacks which target unprotected embeddings [2024:bob], reducing their success rate from 99% to $2\%$. However, as we will show, prior protection schemes are still breakable and do not enjoy cryptographically strong properties that are ideally desired.
Our Pipe attack.
To highlight the gap between the ideal security properties desired and that offered by partially-secure post-processing schemes, we first present a new adaptive attack called Pipe  [^2] Pipe stands for ‘Post-leakage Inversion of Protected Embeddings’. Pipe is a model inversion attack that targets vector embeddings generated using ML models for face images, and post-processed with prior protection mechanisms. Briefly, Pipe recovers a surrogate embedding vector from the protected embedding, then leverages a separate inversion model to recover images close to the original input image. It assumes full leakage—the adversary gets access to all the information stored persistently on a database, including protected embedding vectors. The attack only needs black-box query access to the target ML model and public datasets characterizing the input (face) distribution, but no knowledge of the training dataset of the target model.
We evaluate our Pipe attack against two of the most potent partial protection mechanisms we know of: Facial-Fuzzy Extractor (Facial-FE) [2021:facial-fe] and Multispace Random Projection (MRP) [2007:mrp]. Pipe highlights the weaknesses of these constructions when applied to face authentication under the realistic threat model of full data leakage. Experimentally, Pipe has attack success rate of over 47% for embeddings generated from state-of-the-art models like Facenet [2015:facenet] and ArcFace [2022:arcface] against both of these partial defenses (see Tables III and IV). In most settings, the observed attack success rate is well over 89%.
Figure 1 shows a successfully recovered face from Facial-FE-protected embeddings using Pipe for a quick illustration. It is close to the original image both visually and as per image similarity metrics like LPIPS [2018:LPIPS], showing the privacy-invasive nature of the attack. The normalized embeddings of these images are also close in $\ell_{2}$ distance.
Towards an ideal defense.
Does there exist a cryptographic primitive for $\ell_{2}$-based authentication that would provably defeat all model inversion attacks, regardless of the underlying ML model, with the full-leakage threat model? Three properties would be required by such a primitive: noise tolerance, security under full leakage, and utility. One can think of it as an analogue to the salt-and-hash mechanism used for passwords, except that this primitive must tolerate noise in the input vector. Additionally, it should be hard to recover an input close to the original input even when given all outputs. We introduce this ideal primitive in Section 3.
As a key conceptual contribution (see Theorem 2), we show that the ideal primitive can be achieved if one constructs a secure fuzzy extractor—a cryptographic construct [2004:fe] that predates model inversion attacks by a decade—for the desired $\ell_{2}$ norm vector comparison. Finally, as a technical contribution, we propose a candidate construction called the L2FE-Hash, a lattice-based fuzzy extractor (FE) that enables $\ell_{2}$ norm vector comparison on hashed outputs. We formally show that it is computationally secure under certain assumptions on the input distribution (see Theorem 1). We empirically show that it has reasonable accuracy for face authentication with state-of-the-art ML models. We also confirm that L2FE-Hash resists prominent prior model inversion attacks as well as our adaptive Pipe attack in our experimental evaluation (see Section 8).
To the best of our knowledge, ours is the first work to provide a $\ell_{2}$-based FE with strong practical security and retain reasonable accuracy on real-world face distributions.
Generalizing beyond faces.
While we have chosen the face authentication system as a concrete application, aligning with the original work on model inversion [2015:model-inversion], the threat of inversion attacks on embedding vectors extends well beyond that. All our security properties, attacks, and definitions are generic—they are defined abstractly for models that produce similar high-dimensional embedding vectors for similar inputs. We define closeness with respect to $\ell_{2}$ distance since it is common in face authentication, but one can easily extend our definitions to any distance metric. One can readily consider modalities that rely on $\ell_{2}$ distance, like audio and text [2016:recon-audio, 2022:text-recon-attack] in future work. Likewise, one can consider other security applications such as image similarity checking for copyright infringement or CSAM detection [2021:neural-hash].
Contributions.
To the best of our knowledge, this is the first work that connects two long-standing but disparate lines of research, namely on model inversion attacks and on cryptographic fuzzy extractors. First, we show a new model inversion attack called Pipe which reliably defeats existing post-processing mechanisms, highlighting the lack of practical defenses with desirable cryptographic properties. Second, we describe a candidate fulfilling all ideal properties practically, called L2FE-Hash. We prove that fuzzy extractors are robust against inversion attacks generically, and show that L2FE-Hash is an $\ell_{2}$-norm fuzzy extractor construction that is empirically robust against Pipe, while having a usable accuracy for face authentication.
2 The Problem
Protecting long-term secrets, like authentication credentials, is an important goal in the event of catastrophic data breaches. We explain how this problem manifests in the context of vector embeddings used in modern ML models for face authentication, following the inspiration set in [2015:model-inversion].
2.1 Application: ML Face Authentication
Figure 2: A face authentication system with Enroll and Auth functions. $\mathcal{M}$ represents models that output unprotected embeddings and $\mathcal{M}_{prot}$ represents the models with post-processing protection mechanisms that output protected embeddings.
Billions of users have had password credentials leaked via data breaches, posing serious threats to security and privacy. As authentication services, for example ML-based face authentication like ClearView [2023:clearview-ai] and crypto wallets [2024:biometric-wallet] store credentials (embeddings) derived from face images persistently, they face a similar risk from breach or theft [2024:outabox-breach].
A typical face authentication system employs two phases: enrollment and authentication, denoted by two functions Enroll and Auth respectively (see Figure 2). During enrollment, the user provides an image to the authentication system. The authentication system calls the Enroll function on an image to extract a vector embedding. Specifically, the Enroll function uses an ML model $\mathcal{M}$ often called a “feature extractor” which generates a representative embedding vector $x\in\mathbb{R}^{m}$ for a given input image $I$. It then stores this embedding vector $x$ in the database along with the username. When a user later tries to log into the system, it calls the Auth function using its face image $I^{\prime}$, possibly different from $I$. Auth uses $\mathcal{M}$ to generate an embedding $x^{\prime}$ from the given input $I^{\prime}$, and retrieves $x$ from the database using the username. If $I$ and $I^{\prime}$ are similar, i.e. different images of the same face, $\mathcal{M}$ produces vectors $x$ and $x^{\prime}$ that are close (say in the standard $\ell_{2}$ distance). So, Auth compares the newly computed $x^{\prime}$ with the retrieved $x$ and accepts the authentication request iff the two are close.
In summary, the authentication server stores a high-dimensional embedding vector $x$ in the database. It expects that a legitimate user will have face images that $\mathcal{M}$ maps to vectors close to $x$. It is equally important for security that an illegitimate user will not be able to present an input image at the time of authentication that $\mathcal{M}$ maps to vectors close to $x$. Modern neural networks and generative AI models generate embedding vectors in various practical ways that satisfy both these properties, with acceptably low error rates, and thus are increasingly being used in production systems.
2.2 Inversion of Unprotected ML Embeddings
Consider an attacker who gets unauthorized access to the database storing sensitive embedding vectors persistently. If no protection mechanism was implemented (i.e. $\mathcal{M}$ outputs were stored directly in Figure 2), then the “unprotected” embedding $x$ retains significant information about the original image, and there is a risk that the original image can be inverted from $x$ approximately. Recent prior works have shown that ML models can be trained [2023:stylegan-recontruction-attack, 2023:id3pm-diffusion-inversion-attack, 2024:bob] to perform such inversion for standard face recognition ML systems ($\mathcal{M}$) such as Facenet [2015:facenet] and ArcFace [2022:arcface]. If successful, the inverted image is close enough to the original face image that it passes authentication with $\mathcal{M}$. Prior works have reported attack success rates of over $90\%$ (see Section 6 too). Thus, highly effective inversion attacks on unprotected embeddings are known, posing a serious privacy risk.
2.3 Protecting Embeddings via Post-processing
A good defense mechanism should protect embedding vectors derived from sensitive inputs when they are persistently stored. Ideally, the defense should not be attack-specific and should generically defeat all feasible attacks. Moreover, we are not seeking solutions for securing persistent database storage via additional trust in hardware [2016:sgx-explained] or in third parties to securely store cryptographic keys.
A promising approach to such a defense is via post-processing methods applied directly on embeddings generated from any trained ML model. They can be easily integrated into the face authentication pipeline by replacing $\mathcal{M}$ with $\mathcal{M}_{prot}$ shown in Figure 2. We focus on post-processing protection schemes in the paper since they can be directly implemented on current models without re-training. Such post-processing defenses exist in prior cryptographic literature. We visit two most prominent such mechanisms in Section 4, which our attack will target.
Many prior defenses are not in the post-processing regime. They are methods that work during model training to reduce the amount of sensitive information held by the resulting embeddings [2022:bido, 2023:defense-gan, 2021:mid, 2024:midre]. But such mechanisms have had limited success. Recent works show that the defended model remains susceptible to prior inversion attacks, with $\geq 30\%$ users being re-identified for models that have accuracy $\geq 80\%$ [2022:bido, 2024:midre, 2023:re-thinking]. Therefore, we leave this alternative line of mechanisms out of scope. Besides, they are complementary to post-processing mechanisms.
2.4 Threat Model
We focus on inversion attacks on protected embeddings, i.e., embeddings that have been post-processed before being stored. We assume the full-leakage threat model, much the same as what password hashing is meant to protect against. We assume that all persistently stored metadata is leaked all at once, including any randomness that is strictly necessary for the legitimate authentication to work. Therefore, the adversary’s goal is as follows: Given any leaked face embedding computed as the output of Enroll, along with requisite randomness, reconstruct an image that resembles the original image. Note that ephemeral data such as $(I,I^{\prime},x^{\prime})$ used during Enroll and Auth is not stored persistently and the attacker does not have access to it.
To keep attack assumptions minimal, we work in the black-box threat model. We adapt the black-box model presented in [2015:model-inversion] and assume the attacker only has query (or API) access to the target model $\mathcal{M}_{prot}$ (see Figure 2). The query exposes protected embeddings, instead of classification results as presumed in [2015:model-inversion] for generality, since ML models today are used in non-classification tasks too. Moreover, protection schemes used by the target models are publicly available algorithms known to the attacker.
We assume that the attacker has access to a public dataset of face images, similar to that assumed in the original attack [2015:model-inversion]. Our public dataset does not overlap with the leaked database, and the attacker can query $\mathcal{M}_{prot}$ on this dataset to obtain corresponding embeddings. These assumptions are practical and common to prior attacks [2024:bob, 2021:ked-mi, 2020:gmi]. This is also analogous to how password crackers use real-world password distributions gleaned from leaked password databases [2019:ml-crack-password, 2019:password-crack-survey].
3 An Ideal Post-Processing Primitive
The need for post-processing protection mechanisms has been recognized in prior cryptographic literature [2011:survey-biocrypto, 2018:survey-biocrypto], but their desired security properties have not been carefully established in the context of model inversion. In particular, no secure mechanisms from prior works satisfy the desired properties, as we will experimentally show in Section 6.
1:procedure Enroll($I,\mathcal{M}$)
2:  $x:=\mathcal{M}(I)$ $\triangleright$ $x\in\mathbb{R}^{n}$
3:  $(y,s):=\textsc{Gen}(x)$ $\triangleright$ $y\in\mathbb{R}^{m}$, $s\leftarrow S$
4:  Persistently store $(y,s)$ in Database
5:end procedure
1:procedure Auth($I^{\prime},\mathcal{M}$)
2:  $x^{\prime}:=\mathcal{M}(I^{\prime})$ $\triangleright$ $x^{\prime}\in\mathbb{R}^{n}$
3:  Load $(y,s)$ from Database
4:  $y^{\prime}:=\textsc{Rep}(x^{\prime},s)$
5:  Return $\textsc{Verify}(y^{\prime},y)$
6:end procedure
Algorithm 1 Enrollment and Authentication with ideal primitive.
3.1 Desired Properties
To systematically characterize the problem, we define the properties an ideal protection primitive should satisfy. We do so informally here and provide formal definitions in Appendix A. The ideal primitive contains two subroutines: Gen and Rep. Gen takes input embedding $x$ and outputs a protected embedding $y$ with auxiliary random seed $s$. The function Rep takes some close input $x^{\prime}$ and helper $s$ to produce $y^{\prime}$. An ideal primitive must satisfy $3$ properties:
  1. 1.
    Correctness (Noise tolerance): For close inputs $x$ and $x^{\prime}$, when $x$ is enrolled, $x^{\prime}$ should authenticate successfully with high probability. In other words, for $\textsc{Gen}(x)=(y,s)$, the reproduced embedding $\textsc{Rep}(x^{\prime},s)=y^{\prime}$ is equal to $y$ with high probability.
  2. 2.
    Security (Fuzzy One-wayness): Given outputs $(y,s):=\textsc{Gen}(x)$, finding any input $x^{\prime}$ close to $x$ is hard.
  3. 3.
    Utility: The entropy of the output embeddings $y$, conditioned on helper $s$, should be comparable to that of the input embeddings $x$.
Authentication based on the ideal primitive.
If such a primitive exists, Algo. 1 shows how it readily works with a face authentication system. The Enroll function receives an enrollment image, applies Gen to the image’s embedding vector $x$, then stores the corresponding protected embedding $y$ in the database persistently. The Auth function receives an authentication image, and checks if it has matching identity with the enrollment image. Specifically, it uses Rep to compute some $y^{\prime}$ using $x^{\prime}$. The Verify function compares $y^{\prime}$ with the stored $y$, and outputs Accept only if $y^{\prime}=y$. The user is allowed to log in only if Verify returns Accept.
TABLE I: Symbols used in this paper with definitions
Symbol Definition
$\mathcal{M}$ Target (feature extractor) model
$\mathcal{M}_{prot}$ Target model with post-processing protection
II Input image obtained during enrollment
$I^{\prime}$ Input image obtained during authentication
$x$ Input vector computed during enrollment
$x^{\prime}$ Input vector computed during authentication
$x^{*}$ Surrogate vector recovered from $y$
$y$ Output protected vector obtained from $x$ (persistent)
$y^{\prime}$ Output protected vector obtained from $x^{\prime}$
$m$ Size of vector $x$
$n$ Size of vector $y$
$l$ Size of ephemeral vector $b$ in L2FE-Hash (Algo. 3)
$t$ Closeness $\ell_{2}$-threshold between $x$ and $x^{\prime}$
$T$ Closeness $\ell_{2}$-threshold between $y$ and $y^{\prime}$
$s$ Random seed vector (persistent)
$\lambda$ Security parameter which controls size of seed $s$
$\zeta$ Entropy of input distribution
Why these properties?
The noise-tolerance Item 1 ensures that protected embeddings can be readily used in security applications that demand $\ell_{2}$-based closeness comparison. For example, in a face authentication system, $\mathcal{M}$ generates unprotected embeddings close in $\ell_{2}$ norm given similar face images. After applying the protection scheme, these embeddings should map to the same protected vector allowing Auth to perform exact comparison with high accuracy.
The recovery of an unprotected embedding close to the original poses a high privacy risk, as prior work has shown that unprotected embeddings can be inverted back into original images. Item 2 ensures that the embedding $x$ is hard to recover, even approximately within a distance $t$, even if we leak the auxiliary seed $s$ and the stored output $y$. So, Item 2 generically defeats inversion attacks under the full-leakage threat model—the protected embedding cryptographically hides the unprotected embedding it is generated from.
Item 3 is necessary to prevent trivial functions that compress inputs into a very small output domain. If output entropy is too small, the adversary would be able to guess inputs by brute-forcing. Moreover, it would render the primitive useless in applications that want to use the output to distinguish various inputs. For example, a primitive designed for face authentication should map input embeddings corresponding to different users to different protected embeddings. Hence, the output entropy required for Item 3 depends on the application use case and is often related to the input entropy. We use the standard computational HILL entropy [1999:extractor-universal-hash-hill] as a measure of the output entropy.
3.2 Relation to Fuzzy Extractors
We will later show formally that the desired properties are achieved by a fuzzy extractor. Briefly, the Gen and Rep subroutines of a fuzzy extractor work similarly to the Gen and Rep subroutines of our ideal primitive, albeit with different security formulations. However, existing fuzzy extractor constructions are designed for $\ell_{0}$ distances, such as the Hamming distance or set difference, to suit the use case of iris or fingerprint authentication [2004:fe]. There is no secure construction of an $\ell_{2}$-based fuzzy extractor, as needed for practical use with face distributions yet. We address this gap by proposing an ideal primitive for an $\ell_{2}$-based authentication system, and give a working candidate construction L2FE-Hash of the ideal primitive in Section 7.
4 Existing Protection Schemes
As a starting point for post-processing defenses, one can consider two cryptographic schemes that are state-of-the-art in their respective family of concepts: a Facial-Fuzzy Extractor and Multispace Random Projection. We detail their constructions and explain why they are only partially secure under the threat model of full leakage. They serve as two main targets for our new attack Pipe presented in Section 5, and experimentally confirm it in Section 6.
4.1 Existing Scheme via Fuzzy Extractor
A fuzzy extractor (FE) is an existing cryptographic primitive that derives high-entropy cryptographic keys from a noisy input source, such as human biometrics [2004:fe]. Specifically, it extracts a uniformly random string from an input vector, such that the same string can be reliably reproduced from other close input vectors. This “error-correcting” property makes FE well-suited for noisy-source authentication.
Background on FE.
A FE primitive is defined by two functions Gen and Rep. The function Gen takes input $x$ and outputs a secret extracted key $r$ and a public helper string $p$. The function Rep takes some close input $x^{\prime}$ and helper $p$ to reproduce $r$. Typically, an FE scheme is instantiated using another cryptographic primitive called a secure sketch, which provides the error-correction property, and a randomness extractor, which ensures uniformity of the secret $r$ [2004:fe]. The security property of FE guarantees that the helper $p$ can be released publicly without revealing any information about $r$, enabling $r$ to be used as a secret key for any downstream cryptographic applications.
FE-based Authentication.
The roles of the FE Gen outputs $r$ and $p$ as secrets and helpers respectively correspond to the outputs $y$ and $s$ of the ideal primitive we defined earlier. Hence, an FE scheme can be easily integrated into the face authentication pipeline in Algo. 1 by substituting with FE’s Gen and Rep functions  [^3] Other methods of integrating FE into face authentication have also been proposed, such as [2017:fe-pke] using the FE secret output $r$ to generate keys for a public-key encryption scheme, though this distinction is not relevant.. Note that only $(r,p)$ must be persistently stored for subsequent use at authentication time.
No secure FE construction prior to our work that performs $\ell_{2}$ norm error-correction is known to the best of our knowledge. To illustrate the lack of such a primitive, we will take the closest FE scheme we know of that approximates for $\ell_{2}$-correction, called Facial-Fuzzy Extractor (or Facial-FE) [2021:facial-fe], and show that it is insecure.
Facial-FE.
Facial-FE claims to provide the only computationally secure FE that works for an approximation of the $\ell_{2}$ metric using a specific lattice-based metric [2021:facial-fe]. It uses dense and efficiently decodable lattices, where each input is decoded to a lattice point, and the difference vector is the helper string. The hashed input and the public helper string are stored as the persistent secret to help reconstruct the original input from the authentication input. The Facial-FE construction from [2021:facial-fe] using the E8-lattice decoder as the Decode sub-algorithm is reproduced as-is in Algo. 2, adapted to use symbols consistent with our work. $H_{k}$ denotes a cryptographic hash with a random key $k$.
Facial-FE is not ideal.
Facial-FE can satisfy Item 1 because the linearity of lattices allows for correction of errors within the Voronoi cell, which is an approximation for correction of errors within some $\ell_{2}$ distance. However, it does not satisfy the security property Item 2. Notably, it releases a public helper string $p$, which contains a string $ss$ (called a secure sketch output). The key issue is that the construction uses a deterministic decoder, i.e., there is no randomization in producing $ss$. Since $c$ is determined by $x$ for a fixed lattice, $ss$ = cxc-x can reveal a lot of information about in $x$. Our attacks in Section 5 take advantage of this weakness and use $ss$ to find an $x^{\prime}$ close to $x$.
1:procedure Gen
2:  Input: $x\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}X$ $\triangleright$ $x\in\mathbb{R}^{m}$
3:  $k\leftarrow\{0,1\}^{\lambda-m}$
4:  $c:=\texttt{Decode}(x)$ $\triangleright$ $c\in\mathbb{Z}^{m}$
5:  ss:=cxss:=c-x
6:  $p:=(ss,k)$ and $r:=H_{k}(x,ss)$ $\triangleright$ $p\in\mathbb{Z}^{\lambda}$
7:  return $(r,p)$ $\triangleright$ $r,p$ are stored persistently
8:end procedure
1:procedure Rep
2:  Input: $(x^{\prime},p)$
3:  Parse $p$ as $(ss,k)$
4:  $c^{\prime}:=x^{\prime}+ss$
5:  $c^{*}:=\texttt{Decode}(c^{\prime})$ $\triangleright$ Decode is E8-decoder
6:  $\tilde{x}:=c^{*}-ss$
7:  return $H_{k}(\tilde{x},ss)$
8:end procedure
Algorithm 2 Facial-FE construction from [2021:facial-fe] Algo. 1, 2, 5, 6.
4.2 Existing Scheme via Random Projection
Another popularly used primitive for protecting embeddings via post-processing is multispace random projection (MRP) [2007:mrp]. There are many projection-based schemes used to protect biometrics, but they are variants of the common MRP building block [2004:biohashing, 2018:iom-hash]. MRP works by projecting input vectors onto a random subspace of lower dimensionality, determined by a user-specific random seed. Then it can perform vector comparisons for authentication in the projected vector space without storing the original vectors.
Background on MRP.
MRP is a randomized technique that projects the input vector along random vectors, first proposed by [2007:mrp]. MRP samples a random Gaussian matrix $R\in\mathbb{R}^{n\times m}$ where each entry is from the standard normal distribution with mean 0 and standard deviation $1$. Each row vector in $R$ thus has a random vector direction in the Euclidean space. It projects the input vector $x\in\mathbb{R}^{m}$ to output vector $y\in\mathbb{R}^{n}$, where $n<m$, as follows:
$y\coloneqq f(x)=\frac{1}{\sqrt{n}}Rx$ (1)
The coordinates of $y$ are thus the projections of $x$ along the row vectors in $R$. MRP assumes that $R$ is stored securely in a database and never leaked. We refer to this threat model as partial leakage. In typical breaches, however, everything stored on the server or device is leaked. The security offered by MRP schemes is not formally defined or critically analyzed in the full-leakage threat model so far.
MRP-based Authentication.
One can use MRP in Algo. 1 by setting $s=R$ and $y$ as shown in Eqn 1. Instead of using exact matching as the Verify function, MRP-based face authentication uses closeness matching, where Verify will accept as long as $(y,y^{\prime})$ are close within some threshold distance. Thus, if the unprotected embedding $x$ at the time of enrollment is close to the $x^{\prime}$ presented during authentication within distance $t$, then the corresponding protected embeddings $y$ and $y^{\prime}$ should be close to each other within distance $T$. If so, then the MRP scheme serves the desired functionality of correctness. Note that both $y$ and the random matrix $R$ are stored persistently in the database.
MRP is not ideal.
MRP is not an ideal primitive because it does not satisfy Item 2. It only satisfies the much weaker non-invertibility property, stated informally below:
Given $s$ and $y$, it is hard to find $x$ exactly.
MRP is non-invertible because $Rx=y$ gives us a set of under-determined linear equations. However, non-invertibility does not preclude approximate recovery, i.e., finding an input vector $x^{\prime}$ within distance $t$ from $x$. Thus, it is a weaker security property than Item 2. We show attacks against MRP in Section 5 that exploit this distinction.
MRP does meet $\ell_{2}$ norm correctness Item 1 because of the phenomenon captured by the Johnson-Lindenstrauss (JL) Lemma [1984:JLLemma, 2003:JLLemma-proof]. The JL Lemma 1 given below ensures that the pairwise distance between all points in a set of $k$ points is maintained within an arbitrarily small factor by random projection maps. In other words, the projected data still retains enough information for authentication to work.
Lemma 1 (Johnson–Lindenstrauss Lemma [2003:JLLemma-proof] rephrased).
For any $0<\epsilon<1$ and any integer $k>0$, let $n$ be a positive integer such that $n\geq\frac{4\ln k}{\epsilon^{2}/2-\epsilon^{3}/3}$. Then for any set $S$ of $k=|S|$ data points in $\mathbb{R}^{m}$ and the randomized map $f\colon\mathbb{R}^{m}\to\mathbb{R}^{n}$ given in  Eqn 1, with high probability for all $x_{i},x_{j}\in S$:
$(1-\epsilon)\|x_{i}-x_{j}\|_{2}^{2}\leq\|f(x_{i})-f(x_{j})\|_{2}^{2}\leq(1+\epsilon)\|x_{i}-x_{j}\|_{2}^{2}$
The JL Lemma tells us how, with an appropriate choice of $n$, MRP can satisfy correctness (noise-tolerance) Item 1. Specifically from the lemma, it can be seen that if the dimensions of the projected space $n=O(\ln k/\epsilon^{2})$, then the distance between any two vectors out of $k$ is maintained with a reasonably small $\epsilon$ multiplicative factor of the original distance in $n$ dimensions. This can be used to select the distance threshold $T$ to use for verification, as it is a function of projected dimensionality $n$, the noise tolerance (distance) of unprotected embeddings $t$, and the total number of users $k$. The tunable constant $\epsilon$ determines $n$ and reliability $\delta$.
We conjecture that an MRP can also satisfy Item 3 since random projections largely transfer the entropy of the input to the output—they do not create too many collisions.
5 The Pipe Attack
We propose a new model inversion attack called Pipe to highlight the gap between existing candidates available from literature and the ideal defense. Our evaluation using Pipe against existing schemes is in Section 6. It confirms the gap experimentally. We will also evaluate newly proposed L2FE-Hash in Section 8 to show the security improvement.
5.1 Attack Setup
As discussed in Section 2, the attacker has black-box query (API) access to the target model $\mathcal{M}$ and a public face dataset. It knows the protection mechanism used by $\mathcal{M}$. No ephemeral secrets or intermediate values computed in Algo. 1 are persistently stored or leaked to the adversary; only the protected embedding vectors $y$ and the corresponding seed $s$ are. For FE-based schemes, the protected embedding $y$ is the extracted key $r$ and the seed $s$ is the public helper $p$. For MRP, the protected embedding $y$ is the projected vector $y$ and the seed $s$ is the projection matrix $R$. All persistently stored data is leaked to the adversary.
Choice of seed.
We consider the attack setup where the seed $s$ is a different value per-user. The use of different seeds is similar to how passwords of each user are hashed with different salt values, which in the context of password hashing, is expected to increase the attacker’s effort per user. It has been considered in prior works on protecting embeddings as well. For example, the MRP seed is referred to as a “user-specific PRN” [2007:mrp], and the Facial-FE seed includes a user-specific deterministic vector $ss$. Choosing the same seed for all users would make our attack much more effective, though that would be an artificial choice.
5.2 Attack algorithm
There are two parts in the Pipe attack. Step 1 is to recover a surrogate embedding from all available information. This step demonstrates the weakness of current protection schemes in the full-leakage setting. Step 2 is to perform inversion on the surrogate embedding to obtain the original image. We train a diffusion model conditioned on these surrogate embeddings to reconstruct original images.
5.2.1 Step 1: Recovering surrogate embedding
The step aims to recover a surrogate vector $x^{*}$ from $y$ and $s$, such that $x^{*}$ contains information about the original $x$ which was used to generate $y$. This step is specific to the protection scheme since the attacker can utilize knowledge about the scheme to devise the attack.
Pipe Step 1 for Facial-FE.
Facial-FE outputs $(p$, $r)$ where $p=(ss,k)$ and $r=H_{k}(x,ss)$, see Algo. 2. Since $r$ is a random value, we attack only using the public $ss$ which persists in the database. Recall from Section 4.1 that ss=cxss=c-x and both $c$ and $x$ are determined by $x$ due to the use of a deterministic lattice. We exploit this weakness. Specifically, we observe that decoded codeword $c$ is a deterministic sparse vector for each $x$, making $ss$ largely resemble the input $x$. Hence the surrogate vector $x^{*}=ss$. Our successful attack implies that Facial-FE does not even satisfy standard security properties of fuzzy extractors.
Pipe Step 1 for MRP.
In the case of MRP, the attacker wants to recover some $x^{*}$ from $y=Rx$. Since $R\in\mathbb{R}^{n\times m}$ and $n<m$, $R$ is a compression map. Solving $x=R^{-1}y$ is an under-determined system of linear equations, so Gaussian elimination is not usable directly. To resolve this problem, we propose to compute a Moore-Penrose inverse [1955:moore-penrose-inverse] or simply, pseudo-inverse of $R$, denoted as $R^{\dagger}\in\mathbb{R}^{m\times n}$ using the singular values of $R$. The matrix $R^{\dagger}$ thus obtained is a right pseudo-inverse of $R$ as $m<n$. By left multiplying $R^{\dagger}$ to the leaked vector $y$, we get an $m$-dimensional vector $x^{*}$.
$R^{\dagger}y=R^{\dagger}Rx=x^{*}$
Since we are using a right pseudo-inverse for left matrix multiplication, we do not have the guarantee that $x^{*}$ is equal to or close to $x$. However, we expect that $x^{*}$ contains information similar to the original $x$ since the pseudo-inverse uses the $n$ singular values of $R$, carrying most of the information about how $R$ compresses $x\in\mathbb{R}^{m}$ to $y\in\mathbb{R}^{n}$.
5.2.2 Step 2: Inverting embedding
Next, we want to invert $x^{*}$ obtained from the previous step to get an image close to the original image $I$. It is generally difficult to invert $\mathcal{M}$ just by knowing a random value $y$, but thanks to Step 1, we have recovered $x^{*}$ that is significantly correlated to the input $x$. Step 2 uses $x^{*}$ as an input to an inversion model to recover $I^{\prime}$.
There are many existing ML models that can be considered for use as inversion models. Most create good quality images but dissimilar images [2023:stylegan-recontruction-attack], or poor quality but representative images [2022:cnn-recontruction-attack, 2024:bob]. Moreover, these models have only been examined in the setting of image inversion from “unprotected” embeddings $x$ and not from approximately recovered $x^{*}$ as needed in our attack. For our inversion model, we adapt the classifier-guided diffusion model by [2021:conditionedDDPM] and use surrogate vectors $x^{*}$ obtained in Step 1 instead of class labels as the guidance vector. The public face datasets are used to train the model to learn an association between the distribution of face images $I$ and surrogate embeddings $x^{*}$. Once trained, the inversion model is fixed for use in attacks. At attack time (post-training), the inversion model then generates images by conditioning on the surrogate $x^{*}$ recovered by Step 1 from the leaked embedding $x$.
Defining Attack Success.
[…]
$\frac{\#(\mathcal{M}(\text{inverted image})\approx_{t}\mathcal{M}(\text{original image}))}{\#\text{test images}}$
6 Experimental Evaluation: Existing Defenses
We evaluate partial protection schemes applied to state-of-the-art feature extractor models. Section 6.2 reports the effectiveness of Pipe and contrasts it to that of random guessing. Section 6.3 reports on the potency of prior attacks.
TABLE II: The table provides details for the chosen feature extractor models. The threshold ($t$), True positive rate (TPR), and false positive rate (FPR) are computed over 1000 same and 1000 different identity pairs from Celeb-A dataset.
Model Pre-processor Input image size Output vector size Distance metric Threshold ($t$) TPR % FPR %
Facenet MTCNN 160x160 512 Euclidean normalised 1.103 89.5 0.8
ArcFace MTCNN 112x112 512 Euclidean normalised 1.101 79.3 7.4
6.1 Methodology & Attack Setup
Models.
[…]
Training of Inversion Models.
Recall that Pipe uses an embedding conditioned diffusion model adapted from the implementation by [2021:conditionedDDPM] as the inversion model in Pipe attack. We use hyperparameters and image size of $64\times 64$ as recommended in [2021:conditionedDDPM], with a learning rate of 1e41e-4, batch_size =64, ema_rate of $0.9999$ and anneal_steps 25000.
Datasets & Ensuring no overlaps.
To illustrate that our attack can work across multiple datasets and face distributions, we use $3$ image datasets: Casia-Webface [2014:casia-webface], Celeb-A [2015:celeba], LFW [21012:lfw]. We perform cross-dataset evaluation. Specifically, the inversion model used in Pipe is trained only on the Celeb-A dataset, while the ASR is reported for all the $3$ datasets separately. Furthermore, the inversion model is tested on $1000$ different users (images) in all $3$ datasets, but trained only on $9177$ user images in the Celeb-A dataset different from the ones tested on. Thus, there is no overlap in identities used in training set of Pipe and those tested.
Distance Thresholds.
To calculate ASR, we first compute the thresholds $t$ by evaluating the distance between 1000 pairs of images of the same user and 1000 pairs of images of different users, given in Table II. The ASR evaluation is performed by comparing the distance of unprotected embeddings with respect to $t$, rather than comparing protected embeddings. This is a sound evaluation step since all protection schemes MRP and Facial-FE guarantee that close inputs (unprotected) have corresponding outputs be close. Recall that for MRP, the Verify function should output Accept if the distance between protected embeddings is below some threshold, i.e. $\text{d}(y,y^{\prime})\leq T$. Nonetheless, by JL Lemma, this is equivalent to $\text{d}(x,x^{\prime})\leq t$. Similarly, for Facial-FE and L2FE-Hash, the Verify function should output Accept if protected embeddings match exactly for the original and inverted images. This only happens when the unprotected embeddings are close in $\ell_{2}$ norm i.e. $\text{d}(x,x^{\prime})\leq t$.
All experiments are performed on NVIDIA A40 GPUs. The inversion model in Pipe takes around $10$ hours to train for CelebA, requiring $25,000$ anneal steps in diffusion.
Random Guessing.
Before analyzing Pipe results at face value, we need to see the advantage of an attacker who guesses randomly. If there are $1000$ possible identities for an embedding, one might naively expect an ASR for random guessing to be $1/1000=0.1\%$. Practically though, feature extractors on natural distributions have false positives and false negatives, leading to a higher ASR. To evaluate ASR of random guessing, we take 1000 images, each belonging to a different user, and measure how many unprotected embeddings fall within $\ell_{2}$ distance $t$ of the chosen embedding. We repeat this procedure 1000 times for all three datasets as well as the two feature extractors. ASR for random guessing is reported in all cross-dataset experimental Tables III and IV and later in Table X, which are about $1\%$ to $8\%$ on average.
6.2 PIPE Against Existing Protection Schemes
We run Pipe against embeddings protected with Facial-FE and MRP. Our results presented here are for the cross-dataset evaluation while using different seeds per user.
TABLE III: Cross-Dataset ASR (in %) of PIPE attack against Facial-FE-protected embeddings and Success Probability of Random Guessing ($\pm$ standard deviation).Training set is Celeb-A.
Test set Model PIPE % Random %
Celeb-A Facenet 100.0 $1.01\pm 0.52$
LFW Facenet 99.5 $1.25\pm 0.7$
CASIA Facenet 99.6 $1.33\pm 0.71$
Celeb-A ArcFace 95.8 $8.19\pm 9.13$
LFW ArcFace 96.2 $1.46\pm 1.33$
CASIA ArcFace 89 $1.57\pm 2.81$
Pipe Against Facial-FE.
Table III shows that Pipe succeeds with a high ASR. When setting up Facial-FE, we configured its parameters to attain $\leq 70\%$ FNR [^5] TPR is 11-FNR. FNR $\leq 70\%$ implies TPR $30\%$. as specified by [2021:facial-fe]. This requires the face vectors to be scaled down. The scales we obtain for Facenet and Arcface embeddings are 5.7 and 1.39 respectively. At this scale, which is required for reasonable accuracy, we notice that the inputs tend to decode to lattice points $c$ that are sparse 512-dimensional vectors with more than 194 zero elements. Consequently, the helper string ss=cxss=c-x contains significant information about $x$ making Facial-FE insecure. This issue cannot be resolved simply by changing the scale factor. As mentioned in [2021:facial-fe] as well, when we experimentally reduce the scale factor to improve security significantly, the TPR becomes unusably small. Thus, the poor trade-off between security and accuracy appears unavoidable for Facial-FE.
Pipe has over $89\%$ ASR against Facial-FE, the prior scheme that claims to be a $\ell_{2}$-based fuzzy extractor.
TABLE IV: Cross-Dataset ASR (in %) of PIPE in full leakage (PIPE), partial leakage (Partial) and Success Probability of Random Guessing ($\pm$ standard deviation) against MRP-protected embeddings. Training set is Celeb-A.
Test set Model PIPE % Partial % Random %
Celeb-A Facenet 99.4 0.9 $1.01\pm 0.52$
LFW Facenet 99.1 0.5 $1.25\pm 0.7$
CASIA Facenet 99.5 0.6 $1.33\pm 0.71$
Celeb-A ArcFace 64.1 7.0 $8.19\pm 9.13$
LFW ArcFace 50.8 1.4 $1.46\pm 1.33$
CASIA ArcFace 47.1 3.5 $1.57\pm 2.81$
Pipe Against MRP.
Table IV shows that Pipe succeeds with a high ASR of over $99.1\%$ for Facenet models, and over $47.1\%$ for ArcFace models. On the contrary, if we look at the partial leakage setting, Pipe performs poorly consistently with an ASR of less than 7%. This is in accordance with expectations since MRP maps embeddings to a smaller dimension randomly, and that randomness is not known in the case of partial leakage. Even though the Celeb-A arcface reconstruction accuracy is relatively higher, it is within 1 standard deviation of the mean of the random guessing success rate. Hence, Pipe reaffirms that though MRP provably satisfies the correctness item 1 due to JL-Lemma, the privacy loss is severe in the full leakage setting.
All results reported in Table IV are for output dimension $n=128$ for which target models retain good performance. Pipe becomes even more effective for larger $n$ and has an ASR of over $40\%$ for $n\geq 32$. The input dimension is $m=512$, the default for ArcFace and Facenet.
Pipe has ASR consistently above $47.1\%$ against MRP under full leakage. It is significantly higher than the ASR under partial leakage, highlighting the security gap between full and partial leakage settings.
6.3 Prior Attacks vs Partial Protection Schemes
We have shown that Pipe has high ASR against partial protection schemes. Are these partial protection schemes effective against prior attacks to begin with?
To answer this, we evaluate $3$ state-of-the-art inversion attacks on face recognition models: Bob [2024:bob], KED-MI [2021:ked-mi], GMI [2020:gmi]. Bob is the state-of-the-art white-box embedding inversion attack which has reported better ASR than attacks preceding it. KED-MI and GMI are state-of-the-art black-box model inversion attacks [^6] The core of these attacks only requires black-box access to the target model, and a separate inversion model trained on the public dataset.  [2024:midre]. They both use GANs as the inversion model and have the best-known performance among black-box model inversion attacks [2022:bido].
Evaluation Setup.
Note that KED-MI and GMI are technically designed to invert for closed-set face classification systems, i.e., to identify if an embedding corresponds to a particular member of a given set of faces. Closed-set classification differs from embedding inversion in that it requires training an additional classifier to map output embeddings to class labels corresponding to identities in the given closed set. To make these attacks workable for embedding inversion, we augment a classification layer on the target model, so that classifier outputs can be fed to these attacks as is. We use two disjoint subsets of the Celeb-A dataset corresponding to the same 1000 users to train and test the classification layer, and another disjoint subset of images to train the inversion model. The respective classifier accuracy is reported as well [^7] No classifier layer is used for Pipe and Bob and is reported as N/A..
There is a difference in how KED-MI and GMI attacks are evaluated. Since these 2 attacks are designed for a closed-set classification, they have to be tested on users whose identities have been seen during classifier training. The attack is considered successful if reconstructed images are classified by the target model to the correct identity class. On the other hand, Pipe and Bob are tested on unseen people (i.e. an open-set classification), which is more general and more challenging. So, Pipe and Bob are successful iff the $\ell_{2}$ distance between the embeddings of the reconstructed images and corresponding original images is less than the threshold. Hence, the main comparison point to Pipe is Bob, which is also designed for (open-set) embedding inversion.
We evaluate both Facenet and ArcFace as target models under both unprotected and protected settings, using protection schemes: Facial-FE and MRP. Please see Appendix D for more details about the implementations of prior attacks.
TABLE V: ASR (in %) of mentioned inversion attacks against unprotected Facenet and ArcFace embeddings. Inversion models are trained and evaluated on Celeb-A.
Attack Model ASR % Classifier Acc %
PIPE Facenet 99.8 N/A
Bob Facenet 99.4
KED-MI Facenet 0.06 87.50
GMI Facenet 0.02
PIPE ArcFace 97.7 N/A
Bob ArcFace 96.4
KED-MI ArcFace 0.16 90.96
GMI ArcFace 0.08
TABLE VI: ASR (in %) of mentioned inversion attacks against Facial FE-protected Facenet and ArcFace embeddings. Inversion models are trained and evaluated on Celeb-A.
Attack Model ASR % Classifier Acc %
PIPE Facenet 100.0 N/A
Bob Facenet 95.4
KED-MI Facenet 0.02 73.9
GMI Facenet 0.02
PIPE ArcFace 95.8 N/A
Bob ArcFace 84.8
KED-MI ArcFace 0.1 53.89
GMI ArcFace 0.06
TABLE VII: ASR (in %) of mentioned inversion attacks against MRP-protected Facenet and ArcFace embeddings. Inversion models are trained and evaluated on Celeb-A.
Attack Model ASR % Classifier Acc %
PIPE Facenet 99.4 N/A
Bob Facenet 0.2
KED-MI Facenet 0.10 71.34
GMI Facenet 0.04
PIPE ArcFace 64.1 N/A
Bob ArcFace 2.1
KED-MI ArcFace 0.00 69.61
GMI ArcFace 0.06
Results.
Tables V, VI and VII show that Bob is an effective inversion attack on unprotected embeddings, achieving an ASR exceeding $96.4\%$ on both Facenet and ArcFace. Bob still achieves over $84.8\%$ ASR with Facial-FE, underlining the weakness of Facial-FE in that a prior attack designed for unprotected embeddings alone is sufficient to recover private information. However, the performance of Bob falls drastically to $2.1\%$ against MRP, suggesting that MRP works better than Facial-FE against Bob. In contrast, our adaptive Pipe attack remains effective for MRP and Facial-FE. It has ASR over $95\%$ against Facial-FE and over $64\%$ against MRP, indicating that Pipe is more powerful than Bob.
Figure 3: Reconstructed images by Pipe and Bob attacks against different protection schemes for Facenet embeddings. Check mark indicates successful authentication while cross indicates failure.
TABLE VIII: LPIPS metric scores for Bob and Pipe attack
Protection Scheme Facenet ArcFace
Bob Pipe Bob Pipe
Unprotected 0.28 0.21 0.20 0.20
Facial-FE 0.28 0.20 0.21 0.19
MRP 0.26 0.21 0.23 0.25
L2FE-Hash 0.49 0.27 0.38 0.27
TABLE IX: FID metric scores for Bob and Pipe attack
Protection Scheme Facenet ArcFace
Bob Pipe Bob Pipe
Unprotected 58.35 18.73 51.65 18.78
Facial-FE 58.67 17.93 52.51 17.03
MRP 69.56 20.87 69.87 20.20
L2FE-Hash 105.37 21.46 69.71 19.61
Realistic images.
Figure 3 presents the representative images reconstructed by Pipe and Bob. Most Pipe reconstructions are accepted as the same face by the target model, which is not the case for Bob. Moreover, as seen in Tables VIII and IX, Pipe achieves a lower (better) visual similarity LPIPS score [2018:LPIPS] and a lower (better) FID score [2017:fid]. This further confirms that Pipe creates perceptually closer reconstructions to the original images.
KED-MI and GMI attacks obtain less than $0.16\%$ ASR on average over 5 trials, with or without a protection scheme. The weak performance of KED-MI and GMI is expected since these attacks rely solely on classification outputs, which are inherently more restrictive than using the embedding vector itself. We qualitatively inspected over $20$ examples returned by these attacks, and we find that the reconstructions are quite far from the original consistently. Our implementation builds on code provided by [2022:bido].
Pipe attack is highly effective against prior partial defenses, with ASR ranging from $64\%$ to over $99\%$. In contrast, the SOTA attack designed for embedding inversion, i.e. Bob attack, is mitigated by MRP, with ASR dropping from over $98\%$ to below $0.1\%$.
7 Our Proposed Scheme: L2FE-Hash
We now propose a new $\ell_{2}$-based fuzzy extractor called L2FE-Hash that satisfies all requirements of an ideal primitive empirically. We formally characterize in Theorem 1 that L2FE-Hash has provable security when the benign distribution (of say faces) is well-behaved in a precise sense, regardless of what the adversary does. We further prove that FE realizes all $3$ properties of ideal primitive in Theorem 2, even under full leakage. Hence L2FE-Hash can serve as a secure ideal primitive construction when conditions of  Theorem 1 are met. Our experiments in Section 8 confirm that L2FE-Hash defeats prior attacks we evaluated and Pipe.
L2FE-Hash.
The L2FE-Hash scheme presented in Algo. 3 is a randomized lattice-based scheme that achieves $\ell_{2}$ error correction and full-leakage security. Its Gen procedure constructs a random q-ary lattice from $A$ and samples a random vector $b$, such that $Ab$ gives us a random lattice point. It then computes a noisy version of the lattice point $Ab+x$ where the input embedding $x$ is treated as the noise. The ephemeral secret information $b$ is hashed using a cryptographic hash function $H_{k}$. Only the public $p$, the hashed secret $r=H_{k}(b)$, and randomness $(A,k)$ needed for authentication later are stored persistently in the database.
The Rep procedure receives $x^{\prime}$. It de-serializes $p$ to obtain $c=Ab+x$ that was computed during Gen. It implicitly checks if $c-x^{\prime}$ decodes to the correct ephemeral secret value by checking if the hash of the decoded value $H_{k}(b)$ matches what is stored in the database. The Decode procedure is the well-known Babai’s Nearest Plane (BNP) algorithm [1986:BNP], which performs the $\ell_{2}$ error correction.
1:procedure Gen
2:  Input: $x\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}X$ $\triangleright$ $x\in\mathbb{Z}_{q}^{m}$
3:  $A\leftarrow\mathbb{Z}_{q}^{m\times l}$, $b\leftarrow\mathbb{Z}_{q}^{l}$, $k\leftarrow\{0,1\}^{\lambda}$
4:  $d:=(A,Ab+x)$
5:  $p:=(d,k)$ and $r:=H_{k}(b)$
6:  return $(r,p)$ $\triangleright$ $r,p$ are stored persistently
7:end procedure
1:procedure Rep
2:  Input: $(x^{\prime},p)$
3:  Parse $p$ as $(A,c,k)$
4:  $\beta:=c-x^{\prime}$ $\triangleright$ $\beta$ is $Ab+(x-x^{\prime})$
5:  $b:=\texttt{Decode}(A,\beta)$ $\triangleright$ Decode is BNP
6:  return $H_{k}(b)$
7:end procedure
Algorithm 3 The L2FE-Hash construction.
L2FE-Hash is ideal.
L2FE-Hash can be considered a construction of the ideal primitive by renaming the secret $r$ as $y$ and the helper $p$ as $s$. The correctness of L2FE-Hash follows from the fact that during Rep, we get a noisy lattice point $\beta$ where the noise $(x-x^{\prime})$ has very small $\ell_{2}$ norm when $x,x^{\prime}$ are close, i.e, are from the same user. Then BNP correction can successfully recover the lattice point $Ab$, which can be seen empirically in the reported accuracy of Paragraph 8.0.0.5.
Intuitively, the security of L2FE-Hash stems from the following observation: An adversary who does not know an embedding $x^{\prime}$ close to the embedding $x$ of the target user, cannot recover the secret $b$ from the publicly leaked outputs $(A,Ab+x)$ and $H_{k}(b)$. It is worth noting that without the hashing step, this scheme would no longer be secure under the full leakage scenario. Revealing bits of $b$ would leak information about the input $x$. Therefore, hashing the secret is critical for achieving ideal security Item 2.
Formally Characterizing Security.
Constructing a fuzzy extractor for an arbitrary family of distributions $\mathcal{X}$ is known to be statistically impossible [2020:fe-impossible]. Therefore, to characterize the computational hardness of breaking L2FE-Hash, one must specialize to specific distributions. If we strictly confine the input face distribution $\cal{X}$ to be discrete Gaussian or uniform over a small interval, then the hardness of recovering $b$ from $(A,Ab+x)$ is exactly the cryptographic hardness of the LWE problem [2005:lwe, 2013:dmq]. Nonetheless, this simplification is too unrealistic, as real-world face distributions are not known to follow any cryptographically hard distribution or have any closed-form mathematical characterization.
As the next best approach, we characterize the support of the distribution $\mathcal{X}$ as a set of balls bounded in some $R,R]^{m}$ space, since our application aims to correct within some $\ell_{2}$ radius for a set of users. We refer to such distributions with high min-entropy $h$ as well-behaved. Under the assumption that $\mathcal{X}$ is uniform over the assumed support, we prove that L2FE-Hash is a secure FE where the number of secure output bits $\kappa$ is upper bounded by some function of $h$.
Theorem 1 makes the claim precise (see Appendix B for the proof). The L2FE-Hash output $r$ is a $\kappa$-bitstring, but equivalently a vector in $\mathbb{Z}_{q}^{n}$ as needed, where $n=\kappa\log_{q}(2)$.
Theorem 1.
Let $l,m,q,R>0$ be integers such that $R\geq\alpha\sqrt{\frac{m}{2\pi e}}q^{1-l/m}$ for some $\alpha>1$. Let $C_{\epsilon}\subseteq[-R,R]^{m}$ be the union of $\beta$ disjoint $\epsilon$-balls. Let $\mathcal{X}=\mathcal{U}(C_{\epsilon})$ be uniform over $C_{\epsilon}$, with min entropy $h=H_{\infty}(\mathcal{X})$. Let Ext be an average case $(l\log(q),\gamma,\kappa,\epsilon_{\text{FE}})$-strong extractor given by universal hashing, where $\gamma=h+m\log(\alpha)-m\log((2R+1))$.
Then Algo. 3 satisfies the security property of an $(C_{\epsilon},\mathcal{X},\kappa,t,\epsilon_{\text{FE}})$-FE for any $\kappa\leq\gamma-2\log(1/\epsilon_{\text{FE}})+2$, assuming that $H_{k}$ in Algo. 3 is indistinguishable from Ext.
Remark 1. The above theorem offers a formal lens to understand when L2FE-Hash is provably secure. The distributional assumptions stated there are attack agnostic and do not constrain the attacker in any way—they merely characterize properties which, if the benign distribution $\cal{X}$ satisfies, the provable security of L2FE-Hash will hold.
Remark 2. Theorem 1 assumes that input $\mathcal{X}$ is uniform over $C_{\epsilon}$, but one could extend it to many useful well-behaved non-uniform distributions. Theorem 1 proves that if $\mathcal{X}$ is uniform over $C_{\epsilon}$, L2FE-Hash outputs are computationally indistinguishable from uniform, hence satisfying FE security. If $\mathcal{X}$ is a non-uniform but well-behaved distribution with a bounded probability mass in each disjoint $\epsilon$-ball (user identity) of $C_{\epsilon}$, we conjecture that L2FE-Hash output will be a non-uniform distribution but still with sufficient min-entropy. So, no poly-time adversary would be able to distinguish L2FE-Hash outputs from random given polynomially many samples, and the analogous security claim follows.
Choosing Practical Parameters.
Theorem 1 specifies the minimum input entropy needed for good security for an abstract characterization of the distribution $\cal{X}$. But, determining entropy $h$ for real distributions is intractable. So, we derive practical parameters such that the protected embeddings give good accuracy (TPR over $60\%$) and tractable runtime for Gen and Rep. Note that once TPR exceeds well over $50\%$, majority voting over multiple independent face samples can boost it arbitrarily close to $100\%$ [2011:majority-vote]. To achieve accuracy of $1-\delta$, it suffices to have $O(\log(1/\delta))$ face images of the user at authentication time, which can be run in parallel for efficiency. We evaluate L2FE-Hash using the best-known attacks, including prior ones and our Pipe attack on real datasets, to practically assess security.
Fuzzy Extractors and Ideal Defense.
We have shown that L2FE-Hash is a provably secure fuzzy extractor for $\ell_{2}$ distance if the input distribution $\cal{X}$ satisfies certain conditions. Next, we state the desired connection between a fuzzy extractor and an ideal defense primitive as Theorem 2. We defer the formal definition of these two primitives and the ideal primitive security game to Appendix A, and the proof of Theorem 2 to Appendix C due to space limits.
Theorem 2.
Let FE be a ($X,\mathcal{X},\kappa,t,\epsilon_{\text{FE}}$)-fuzzy extractor as in [2004:fe].Then FE is a ($X,\mathcal{X},\{0,1\}^{\kappa},t,\lambda$)-ideal primitive with adversarial advantage $\text{Adv}(\mathcal{A})\leq\epsilon_{\text{FE}}(1-\delta)$.
8 Experimental Evaluation: L2FE-Hash
Following the same methodology and attack setup in Section 6, we now evaluate the security of L2FE-Hash experimentally against Pipe and other prior attacks we tested. We configure Pipe for L2FE-Hash as follows.
Pipe Step 1 for L2FE-Hash.
For L2FE-Hash, similar to Facial-FE, we have public information $p=(A,Ab+x,s)$ and private information $r=H_{s}(b)$. Since $r$ is random, it cannot help the attack in any way. Hence in Step 1, we append the flattened matrix $A$ to $Ab+x$, and use that as $x^{*}$.
Practical L2FE-Hash Parameters.
All evaluations use embedding $x$ projected to $180$ dimensions for computational efficiency. The embedding is quantized [^8] For embedding $x$, the quantized embedding is $\lfloor(x-\min)/(\max-\min)*q*\text{scale}\rfloor$ where min and max are the smallest and largest elements of any embedding vector in the Celeb-A train dataset. to $\mathbb{Z}_{q}$, and then scaled down by a scale hyperparameter optimized for BNP correction. The hyperparameters used for L2FE-Hash protection scheme are $m=180$, $l=60$, $q=130003$ with $\text{scale}=0.0014$, $0.0017$ for Facenet and ArcFace respectively. These remain the same throughout our evaluation.
TABLE X: Cross-Dataset ASR (in %) of PIPE attack against L2FE-Hash-protected embeddings and Success Probability of Random Guessing ($\pm$ standard deviation). Training set is Celeb-A.
Test set Model PIPE % Random %
Celeb-A Facenet 0.6 $1.01\pm 0.52$
LFW Facenet 0.5 $1.25\pm 0.7$
CASIA Facenet 0.4 $1.33\pm 0.71$
Celeb-A ArcFace 7.0 $8.19\pm 9.13$
LFW ArcFace 1.7 $1.46\pm 1.33$
CASIA ArcFace 3.3 $1.57\pm 2.81$
Pipe Against L2FE-Hash.
Table X reports the ASR of Pipe against L2FE-Hash to be at most 7% for all datasets. This is within one standard deviation of the mean ASR for random guessing, which is expected for a secure scheme.
TABLE XI: ASR (in %) of mentioned inversion attacks against L2FE-Hash-protected Facenet and ArcFace embeddings. Inversion models are trained and evaluated on Celeb-A dataset.
Attack Model ASR % Classifier Acc %
PIPE Facenet 0.60 N/A
Bob Facenet 0.60
KED-MI Facenet 0.06 50.33
GMI Facenet 0.04
PIPE ArcFace 7.0 N/A
Bob ArcFace 7.0
KED-MI ArcFace 0.1 38.33
GMI ArcFace 0.08
Prior Attacks against L2FE-Hash.
Table XI shows that all $4$ attacks—Pipe, Bob, GMI and KED-MI—have poor ASR against L2FE-Hash, unlike for partial protection schemes. The ASR is below $0.6\%$ for Facenet embeddings and below $7\%$ for ArcFace embeddings, and always within one standard deviation of the average ASR of random guessing.
All evaluated attacks, including Pipe, fail consistently against L2FE-Hash, highlighting its practical security.
TABLE XII: BNP correction results (in %) against L2FE-Hash embeddings on Celeb-A dataset. The number in $(\cdot)$ indicates the scale by which embeddings are compressed in the $\mathbb{Z}_{q}$ space.
Facenet (0.0014) ArcFace (0.0017)
Actual $\backslash$ Predicted Positive Negative Positive Negative
Positive 65 35 60 40
Negative 4 96 26 74
Accuracy of Face Recognition with L2FE-Hash.
To be practically usable, L2FE-Hash must have reasonably good accuracy, while defending against model inversion. We evaluate L2FE-Hash for 100 pairs of same user images and 100 pairs of different user images from Celeb-A dataset. From Table XII, we observe that L2FE-Hash has above $60\%$ TPR for both target models (feature extractors) with FPR less than $26\%$. Since the TPR $\geq 50\%$ and FPR $\leq 50\%$, one can employ majority voting over multiple independent samples of the user’s face to boost the accuracy overall as needed, as previously mentioned in Section 7 [2011:majority-vote].
9 Related Work
An embedding inversion attack is a canonical example of model inversion attacks that targets model embeddings [2015:model-inversion]. A number of model inversion attacks on image recognition ML models have been developed, often targeting image classification outputs [2021:ked-mi, 2020:gmi, 2021:vmi, 2023:re-thinking, 2023:plgmi]. There are also model inversion attacks on face authentication systems that target image embeddings [2024:bob, 2023:id3pm-diffusion-inversion-attack]. We have evaluated $3$ prior attacks representative of SOTA attack accuracy.
Gradient inversion attacks [2019:gradient-inversion] require more than just model outputs and do not meet the original definitions of model inversion [2015:model-inversion]. They are not in scope of our defenses.
Non-post-processing defenses.
Many defenses against model inversion modify the training process, not post-processing outputs of ML models already trained [2022:bido, 2021:mid, 2024:midre, 2023:defense-gan]. Being highly model-specific and optimization-based, it is hard to theoretically characterize their end security guarantees. Their recently reported effectiveness against existing black-box attacks is fairly limited, with attacks re-identifying $\geq 30\%$ of users from reconstructed images from models that have accuracy over $\geq 80\%$ [2024:midre, 2022:bido]. Importantly, our work points towards defining properties of cryptographic defenses that generically work with ML models for authentication and in an attack-agnostic way.
The privacy risk from model inversion is immediate, which is our primary concern. There is also re-authentication risk, i.e., using recovered face images to impersonate the user. Liveness detectors during face authentication mitigate such risks, but a line of prior works originating with Xu et. al. [2016:defeat-liveness-detect] shows attacks against such mechanisms. These techniques are complementary to mechanisms we study.
Post-processing protection schemes.
Most post-processing protection mechanisms for embedding vectors are some variant of MRP, or use other non-invertible transformations that do not satisfy fuzzy one-wayness (Item 2). For example, BioHashing is MRP followed by a binary quantization step, and it is known to have pre-image attacks [2004:biohashing, 2005:bh-preimage-attack-withR, 2009:bh-preimage-attack-withoutR]. Index-of-Max Hashing performs random projection on fingerprint minutiae and is also vulnerable [2018:iom-hash, 2020:iom-hash-attack]. In fact, many protection mechanisms that rely on a distance-preserving transformation are vulnerable to pre-image attacks [2024:rp-attack, 2019:genetic-preimage-attack], and thus do not satisfy Item 2. There are also schemes that combine MRP with fuzzy commitment for additional security, but they still do not address inversion attacks that recover approximately close inputs [2019:rp-fc, 2021:rp-fc-noninvert].
Similarly, fuzzy hashes, or locality-sensitive hashes [1999:locality-sensitive-hash, 2014:locality-preserve-hash] aim to find linear projections that preserve local information. For example, Apple’s NeuralHash trains a CNN model to learn such mappings for security applications, but it remains vulnerable to adversarial attacks that force hash collisions [2021:neural-hash, 2022:neural-hash-attack]. Furthermore, CNN models in general are particularly susceptible to model inversion attacks [2022:cnn-recontruction-attack].
Other fuzzy extractor schemes have been proposed, such as the LWE-FE [2020:lwe-fe]. While being similar to our L2FE-Hash, it has a different decoder construction that targets the ($\ell_{0}$ norm) Hamming distance. Moreover, it assumes the input distribution to be uniform over a small interval or LWE-admissible, which is an extremely strong assumption for real face distributions. To our knowledge, no prior work gives a secure FE construction for $\ell_{2}$ distances under assumptions that even approximately models real-world distributions and yet shows reasonable accuracy in practical setups.
Other Threat Models.
Our L2FE-Hash security definition under full-leakage threat model is not to be confused with post-compromise security. The latter addresses the issue of secure encryption of future messages after key leakage, whereas Item 2 addresses the issue of information leakage about past messages after key leakage [2016:post-compromise].
Item 2 is also different from forward secrecy, a property of key-exchange protocols ensuring that compromising a long-term key does not reveal past session keys [2019:forward-secrecy]. The fundamental idea that a compromise should not lead to decrypting past messages is similar to L2FE-Hash, but our setting is different from a key-exchange protocol. We additionally assume past encrypted messages, which in our context are protected embeddings, can be compromised.
Honey encryption (HE) addresses a different threat model from ours [2014:honey-enc]. It allows for encryption of passwords with low entropy keys such that decryption with incorrect keys yields honey messages—messages that occur as likely as the original message but are hard to guess without a database breach [2014:honey-enc, 2024:bernoulli-honeywords]. We consider the event of full database breach. HE makes it difficult for the adversary to distinguish correct from incorrect decryption, thwarting offline message recovery attacks. It can aid in breach detection by alerting when an adversary unknowingly queries the database with the honeyword. Moreover, password encryption schemes generally do not require locality-sensitivity.
10 Conclusions
We have shown that fuzzy extractors offer an attack-agnostic and model-agnostic defense against model inversion attacks on embeddings leaked from a database breach. Our Pipe attack demonstrates that prior partially-secure primitives are insecure in this context. We have proposed L2FE-Hash, a fuzzy extractor for vector inputs that enables comparators in Euclidean norm, offering strong security even when all persistently stored data is leaked in the breach.
11 Ethical Considerations
All the datasets used in this paper, Casia-Webface [2014:casia-webface], Celeb-A [2015:celeba] and LFW [21012:lfw], have licenses which allow their use for face verification tasks for research purposes. We do not foresee ethical issues arising from our work, as we have shown both powerful attacks and a strong defense.
Acknowledgment
We would like to thank Divesh Aggarwal for insightful discussions that helped clarify the proof of Theorem 1. This research is supported by a Singapore Ministry of Education (MOE) Tier 2 grant MOE-T2EP20124-0007.
Appendix A Formalization of the Ideal Primitive
We present the formal definition of the idealized primitive introduced in Section 3, with the formal security game.
Notation.
Aside from symbols specified in Table I, we use uppercase alphabets for sets, metric spaces, and random variables. We use uppercase cursive alphabets for probability distributions. Specifically, $\mathcal{U}(S)$ denotes the uniform random distribution over the set $S$. We use $\stackrel{{\scriptstyle\mathcal{D}}}{{\leftarrow}}Q$ to denote sampling from a set $Q$ according to distribution $\mathcal{D}$, or simply $\leftarrow\mathcal{D}$ when the underlying set is obvious from context. $X$ and $Y$ are input and output sets from a metric space with $\text{d}=\ell_{2}$ distance, such as $X=\mathbb{R}^{m}$ and $Y=\mathbb{R}^{n}$. The ball $B_{m}(x,t)$ is the set of all vectors within distance $t$ of $x\in\mathbb{R}^{m}$. We also use $B(x,t)$ or $B_{m}(t)$ to denote the ball $B_{m}(x,t)$ when the omitted variable is clear from context.
A.1 Formal definition
Definition 3 (Ideal primitive).
A ($X,\mathcal{X},Y,t,\lambda$) ideal primitive is a pair of randomized procedures, “generate” (Gen) and “reproduce” (Rep). Gen on input $x\in X$ outputs a seed $s\in\{0,1\}^{\lambda}$ and $y\in Y$. Rep takes $x^{\prime}\in X$ and $s\in\{0,1\}^{\lambda}$ as inputs and gives some $y^{\prime}\in Y$ as output. (Gen, Rep) have the following properties:
  1. 1.
    Correctness (Noise tolerance): If $\text{d}(x,x^{\prime})\leq t$ and $(s,y)=\textsc{Gen}(x)$, then
    ${\Pr}[y=\textsc{Rep}(x^{\prime},s))]\geq 1-\delta.$
  2. 2.
    Security (Fuzzy one-way): Given $(s,y)$ from $\textsc{Gen}(x)$ where $x\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}X$, it is hard to find $x^{\prime}$ such that $\text{d}(x,x^{\prime})\leq~t$. In other words, for any computationally bounded adversary $\mathcal{A}$ trying to find such $x^{\prime}$, its advantage i.e. $\text{Adv}(\mathcal{A})\leq\text{negl}(\lambda)$ (see Eqn 2 for definition of Adv).
  3. 3.
    Utility (Entropy sufficiency): Let $\mathcal{Y}$ and $\mathcal{S}$ be the output distributions of Gen for an input distribution $\mathcal{X}$. The HILL entropy of $\mathcal{Y}|\mathcal{S}$ is at least the min entropy of $\mathcal{X}$.
The probability in Item 1 is taken over $x^{\prime}\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}{B}(x,t)$ and $x\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}X$. The $\delta$ term in Item 1 can be any value in $[0,1]$. It only affects functional correctness, such as the false positives and negatives in the authentication application. On the other hand, “hardness” in Item 2 is cryptographic, i.e., the probability of an adversary violating Item 2 should be a negligible function in $\lambda$. We consider a computationally bounded adversary that runs in $\text{poly}(\lambda)$ time. We formalize the security game of Item 2 in Section A.2. In Item 3, HILL entropy is the computational notion of min entropy [1999:extractor-universal-hash-hill].
A.2 Formal Security game
To make Item 2 precise, we define two notions of indistinguishability below. We use the computational version for Item 2 and the security game is given below.
Definition 4 (Statistical indistinguishability).
Let $X$ and $Y$ be random variables with supports $S_{X}$ and $S_{Y}$ respectively. The statistical distance between X and Y is defined as
$\textbf{SD}(X,Y):=\frac{1}{2}{\sum_{u\in S_{X}\cup S_{Y}}}|\Pr[X=u]-\Pr[Y=u]|$
[…]
Definition 5 (Computational indistinguishability).
Let $X$ and $Y$ be random variables. For a probabilistic polynomial-time (PPT) algorithm $\mathcal{A}$ and a security parameter $\lambda$, the adversarial advantage of distinguishing them is defined as
$\text{Adv}(\mathcal{A})=|\Pr[\mathcal{A}(X)=1]-\Pr[\mathcal{A}(Y)=1]|$
[…]
Security Game.
Let the input distribution $\mathcal{X}$ over $X$ be public information. The game $G$ consists of adversary $\mathcal{A}$ and challenger $\mathcal{C}$ doing the following:
  1. 1.
    Challenger $\mathcal{C}$ samples $x\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}X$, computes $(s,y)=\textsc{Gen}(x)$, and gives $(s,y)$ to adversary $\mathcal{A}$.
  2. 2.
    $\mathcal{A}$ receives $(s,y)$ and outputs $x^{\prime}$.
  3. 3.
    $\mathcal{A}$ wins if $x^{\prime}\in B(x,t)$. $G$ returns 1 if $\mathcal{A}$ wins, else 0.
The advantage of the adversary $\mathcal{A}$ is defined as:
$\text{Adv}(\mathcal{A})=\Pr[G(\mathcal{A},\mathcal{C})=1]-\Pr[G(\mathcal{A}^{\prime},\mathcal{C})=1]$ (2)
where $\mathcal{A}^{\prime}$ is the best baseline adversary that only utilizes $s$ and $\mathcal{X}$ but not $y$ at step 2. If $\text{Adv}(\mathcal{A})$ is negligible for all adversaries $\mathcal{A}$, then Item 2 is satisfied.
Appendix B Theoretical Security of L2FE-Hash
We prove that L2FE-Hash satisfies FE security under the assumption that the input distribution is uniform over a bounded support consisting of multiple disjoint $\epsilon$-radius balls, as in Theorem 1. The center of each $\epsilon$-ball can be seen as the representative embedding of a user. Then the $\epsilon$-ball around the center represents the cluster of embeddings corresponding to that user.
B.1 Definitions
A fuzzy extractor (FE) extracts a uniformly random string with error tolerance. It was first proposed by [2004:fe], and [2020:fe-impossible] relaxed the definition to account for $\delta$ error in correctness. We adapt the definition of FE with $\delta$ errors by [2020:fe-impossible] for a specific input distribution in Definition 6. Let $M$ be a metric space with distance metric d, and $\mathcal{X}$ be a distribution over $M$.
Definition 6 (Fuzzy Extractor with Error).
An ($M,\mathcal{X},\kappa,t,\epsilon_{\text{FE}}$)-fuzzy extractor with error $\delta$ is a pair of randomized procedures, “generate” (Gen) and “reproduce” (Rep). Gen on input $w\in M$ outputs an extracted string $r\in\{0,1\}^{\kappa}$ and a helper string $p\in\{0,1\}^{*}$. Rep takes $w^{\prime}\in M$ and $p\in\{0,1\}^{*}$ as inputs, and outputs $r^{\prime}\in\{0,1\}^{\kappa}$. (Gen, Rep) have the following properties:
  1. 1.
    Correctness: If $\text{d}(w,w^{\prime})\leq t$ and $(r,p)=\textsc{Gen}(w)$, then $\Pr[\textsc{Rep}(w^{\prime},p)=r]\geq 1-\delta$. If $\text{d}(w,w^{\prime})>t$, then no guarantee is provided about the output of Rep.
  2. 2.
    Security: For input distribution $\mathcal{X}$, if $(\mathcal{R},\mathcal{P})=Gen(\mathcal{X})$, then $(\mathcal{R},\mathcal{P})\approx_{\epsilon_{\text{FE}}}(\mathcal{U}(\{0,1\}^{\kappa}),\mathcal{P})$.
The security property of Definition 6 requires the string $r$ to be nearly uniform even for those who observe $p$. The notation $\approx_{\epsilon_{\text{FE}}}$ represents both notions of indistinguishability as defined in Section A.2.
The strong extractor primitive (Definition 7) is commonly used to construct fuzzy extractors. It can be implemented as a universal hash function, and can extract $\kappa\leq\gamma-2\log(1/\epsilon_{\text{FE}})+2$ many bits of randomness, effectively transforming a high entropy input (Definition 10) into a uniformly distributed output (see Lemma 2.4 in [2004:fe]).
Definition 7 (Strong extractor [2004:fe, Section 2.5]).
Let $\text{Ext}\colon\{0,1\}^{l}\to\{0,1\}^{\kappa}$ be a polynomial time probabilistic function using $\lambda$ bits of randomness. Ext is an efficient average-case $(l,\gamma,\kappa,\epsilon_{\text{FE}})$-strong extractor if for all pairs of random variables ($X,I$) such that $X$ is an $l$-bit string satisfying $\tilde{H}_{\infty}(X|I)\geq\gamma$, we have $\textbf{SD}((\text{Ext}(X;K),K,I)$, $(\mathcal{U}(\{0,1\}^{\kappa}),K,I))\leq\epsilon_{\text{FE}}$, where $K$ is uniform on $\{0,1\}^{\lambda}$.
Now we present the definition of several terms used to characterize the input distribution and used in the proof.
Definition 8 ($C_{\epsilon}$).
For $\epsilon>0$ and some integer $R>0$, let $C\subset[-R+\epsilon,R-\epsilon]^{m}$ be a closed set of size $\beta$ such that $\forall c_{1},c_{2}\in C$ where $c_{1}\neq c_{2}$, $\|c_{1}-c_{2}\|_{2}>2\epsilon$. Then $C_{\epsilon}$ denotes the union of disjoint $\epsilon$-balls with centers in $C$, i.e.
[…]
Definition 9 (Ambiguity).
We say that a vector $\mathbf{y}\in\mathbb{Z}_{q}^{m}$ is $N$-ambiguous for a matrix $\mathbf{A}$ and an error distribution $\mathcal{X}$, if $|\{\mathbf{b}\in\mathbb{Z}_{q}^{l}\mid\mathbf{y}-\mathbf{A}\mathbf{b}\in\text{supp}(\mathcal{X})\}|\geq N$. If $A$ and $\mathcal{X}$ are clear by context, we just say that $\mathbf{y}$ is $N$-ambiguous.
Definition 10 (Min Entropy [2004:fe, Section 2.4]).
Let log denote the base 2 logarithm. Let $(X,Y)$ be a pair of random variables. Define min entropy of $X$ as $H_{\infty}(X)=-\log(\max_{x}\Pr[X=x])$, and the average (conditional) min-entropy of X given Y as
$\tilde{H}_{\infty}(X|Y)=-\log\left(\mathbb{E}_{y\leftarrow Y}\max_{x}\Pr[X=x|Y=y]\right).$
B.2 Proof of Intermediate Theorem
Before proving Theorem 1, we first prove that the public information $d=(A,Ab+x)$ of L2FE-Hash in Algo. 3 has high entropy, as stated in Theorem 11.
Theorem 11.
Let $l,m,q,R>0$ be integers such that $R\geq\alpha\sqrt{\frac{m}{2\pi e}}q^{1-l/m}$ for some $\alpha>1$. Let $\mathcal{X}=\mathcal{U}(C_{\epsilon})$ and $h=H_{\infty}(\mathcal{X})$. Let $\mathbf{A}$ be distributed according to $\mathcal{U}(\mathbb{Z}_{q}^{m\times l})$. Let $\mathbf{y}=\mathbf{A}\mathbf{b}^{\prime}+\mathbf{x}^{\prime}$, with $\mathbf{b}^{\prime}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{l})$ and $\mathbf{x}^{\prime}\leftarrow\mathcal{X}$. Also let $\mathbf{b}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{l})$ and $\mathbf{x}\leftarrow\mathcal{X}$, then with probability $1-1/q^{m-l}$,
$\tilde{H}_{\infty}(\mathbf{b}|\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y})\geq\gamma(h)$
where $\gamma(h)=h+m\log(\alpha)-m\log((2R+1))$.
Proof of Theorem 11.
Define $p$ to be the probability that an integer point chosen uniformly at random from $R,R]^{m}$ lands in $C_{\epsilon}$. We get Fact 1: that $p:=\frac{f(\beta,\epsilon,m)}{(2R+1)^{m}}=\frac{2^{h}}{(2R+1)^{m}}$ since by Lemma 3, we know that $f(\beta,\epsilon,m)=2^{h}$.
Consider a random matrix $\mathbf{A}\in\mathbb{Z}_{q}^{m\times l}$. The corresponding q-ary lattice is $\mathcal{L}(\mathbf{A})=\{x\mid x=\mathbf{A}s\mod q\text{ for some }s\in\mathbb{Z}_{q}^{l}\}\subseteq\mathbb{Z}_{q}^{m}$. Then the expected number of lattice points inside $C_{\epsilon}$ is
${\mathbb{E}_{\mathbf{A}}[\#(\mathcal{L}\cap C_{\epsilon})]}=~{\mathbb{E}_{\mathbf{A}}[\sum_{v\in\mathcal{L}}1(v\in~C_{\epsilon})]}\\ =~{\sum_{v\in\mathcal{L}}\mathbb{E}_{\mathbf{A}}[1(v\in C_{\epsilon})]}=~{\sum_{v\in\mathcal{L}}\Pr[v\in C_{\epsilon}]}\\ =~{\sum_{v\in\mathcal{L}\cap[-R,R]^{m}}{\Pr[v\in C_{\epsilon}]}}=p\cdot\#(\mathcal{L}\cap[-R,R]^{m})$ where the fourth equality is by $\Pr[v\in C_{\epsilon}]=0$ for $v\not\in[-R,R]^{m}$.
By Minkowski’s Convex Body Theorem, $\frac{\text{vol}(S)}{\det(\mathcal{L})}$ estimates the number of lattice points in set $S$ [1996:minkowski-textbook]. Let $S$ be $B_{m}(R)$, the $m$-dim. sphere of radius $R$ inscribed within the $R,R]^{m}$. Then for a random lattice $\mathcal{L}(\mathbf{A})$ where $\mathbf{A}$ is full rank, the number of lattice points in $B_{m}(R)$ is
$\begin{split}&\mathbb{E}[\#(\mathcal{L}\cap B_{m}(R))]=\frac{\text{vol}(B_{m}(R))}{\det(\mathcal{L})}=R^{m}\frac{\text{vol}(B_{m}(1))}{\det(\mathcal{L})}\\ &=R^{m}\frac{\text{vol}(B_{m}(1))}{q^{m-l}}=R^{m}\frac{\pi^{m/2}/(m/2)!}{q^{m-l}}\approx R^{m}\frac{(2\pi e/m)^{m/2}}{q^{m-l}}.\end{split}$ (3)
The third equality holds because $\det(\mathcal{L}(\mathbf{A}))=q^{m-l}$ for a full-rank matrix $\mathbf{A}$, and the last $\approx$ is by Stirling’s approximation. By Lemma 4.4 of [2020:lwe-fe], $\mathbf{A}$ is full rank with probability at least $1-1/q^{m-l}$ as needed in the Theorem.
Define $\lambda_{1}$ to be the largest radius such that $\mathbb{E}[\#(\mathcal{L}\cap B_{m}(\lambda_{1}))]=1$. If we set $R=\lambda_{1}$ in Eqn 3, we can get the expression $\mathbb{E}[\lambda_{1}]=\sqrt{\frac{m}{2\pi e}}q^{1-l/m}$. Now consider setting $R\geq\alpha\mathbb{E}[\lambda_{1}]$. By Eqn 3, $\mathbb{E}[\#(\mathcal{L}\cap B_{m}(R))]\geq\mathbb{E}[\#(\mathcal{L},B_{m}(\alpha\mathbb{E}[\lambda_{1}]))]\allowbreak=(\alpha\mathbb{E}[\lambda_{1}])^{m}\frac{(2\pi e/m)^{m/2}}{q^{m-l}}\allowbreak=\alpha^{m}$. Then for $C_{\epsilon}\subset[-R,R]^{m}$ (see Definition 8), we get
$\mathbb{E}[\#(\mathcal{L}\cap~C_{\epsilon})]\geq p\cdot\alpha^{m}=\frac{2^{h+m\log(\alpha)}}{(2R+1)^{m}}$ (4)
where the inequality and equality are directly by Fact 1.
We take log on both sides of Eqn 4, and set $\gamma=\log\left(\frac{2^{h+m\log(\alpha)}}{(2R+1)^{m}}\right)=h+m\log(\alpha)-m\log(2R+1)$ as in Theorem statement. So, we get Fact 2: $\mathbb{E}_{\mathbf{A}}[\#(\mathcal{L}\cap C_{\epsilon})]\geq 2^{\gamma}$.
[…]
$\mathbb{E}_{\mathbf{A}}\max_{\mathbf{b}}\Pr[\mathbf{b}|\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y}]\leq 2^{-\gamma}.$ (5)
Since $\mathbf{y}$ is arbitrary, Eqn 5 holds for all $\mathbf{y}$. By Definition 10, $\tilde{H}_{\infty}(\mathbf{b}|\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y})\allowbreak=-\log(\mathbb{E}_{\mathbf{A},\mathbf{y}}\max_{\mathbf{b}}\Pr[\mathbf{b}|\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y}])\allowbreak\geq-\log(2^{-\gamma})=\gamma.$
Lemma 2.
Let $\beta\in\mathbb{Z},m\in 2\mathbb{Z},\epsilon\in\mathbb{R}>0$. Then $C_{\epsilon}\subseteq[-R,R]^{m}$ satisfies $f(\beta,\epsilon,m):=|C_{\epsilon}|=\beta\cdot\frac{\pi^{m/2}}{(m/2)!}\epsilon^{m}$.
Proof.
The volume of $B_{m}(\epsilon)$ is $\text{vol}(B_{m}(\epsilon))=\frac{\pi^{m/2}}{\Gamma(m/2+1)}\epsilon^{m}$, or simply $\text{vol}(B_{m}(\epsilon))=\frac{\pi^{m/2}}{(m/2)!}\epsilon^{m}$ for even mm [^9] As $m$ is typically even in practice, we assume so henceforth.. Then $\begin{aligned} |C_{\epsilon}|=\beta\cdot\text{vol}(B_{m}(\epsilon))=\beta\cdot\frac{\pi^{m/2}}{(m/2)!}\epsilon^{m}\end{aligned}$ by Definition 8. ∎
Lemma 3.
For $\mathcal{X}=\mathcal{U}(C_{\epsilon})$, $H_{\infty}(\mathcal{X})=\log(f(\beta,\epsilon,m))$.
Proof.
By Definition 10, $H_{\infty}(\mathcal{X})=H_{\infty}(\mathcal{U}(C_{\epsilon}))=-\log(\max\Pr[\mathcal{U}(C_{\epsilon})])=-\log\left(\frac{1}{|C_{\epsilon}|}\right)$. By Lemma 2 $H_{\infty}(\mathcal{X})=\frac{1}{f(\beta,\epsilon,m)}=\log(f(\beta,\epsilon,m)).$
Lemma 4.
Fix a matrix $\mathbf{A}\in\mathbb{Z}_{q}^{m\times l}$. Let $\mathbf{y}\in\mathbb{Z}_{q}^{m}$ be $2^{\gamma}$-ambiguous for the matrix $\mathbf{A}$ and error distribution $\mathcal{X}=\mathcal{U}(C_{\epsilon})$. Let $\mathbf{b}\leftarrow\mathcal{U}(\mathbb{Z}_{q}^{l})$ and $\mathbf{x}\leftarrow\mathcal{X}$. Then for all $\mathbf{z}\in\mathbb{Z}_{q}^{l}$,
$\Pr[\mathbf{b}=\mathbf{z}|\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y}]\leq 2^{-\gamma}.$
In other words, $\max_{\mathbf{b}}\Pr[\mathbf{b}|\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y}]\leq 2^{-\gamma}$.
The proof is similar to that of Lemma 4 in [2013:dmq], since $\mathcal{X}$ is also a uniform distribution over its support.
Proof.
Since $\mathbf{b}$ and $\mathbf{x}$ are drawn from uniform distributions, $p:=Pr[\mathbf{b}=\tilde{\mathbf{b}},\mathbf{x}=\tilde{\mathbf{x}}$] is the same for all $\tilde{\mathbf{b}}\in\mathbb{Z}_{q}^{l}$ and $\tilde{\mathbf{x}}\in\mathcal{X}$. Let $S:=\{\mathbf{z}\in\mathbb{Z}_{q}^{l}|\mathbf{y}-\mathbf{A}\mathbf{z}\in\text{supp}(\mathcal{X})\}$. As $\mathbf{y}$ is $2^{\gamma}$-ambiguous it holds that $|S|\geq 2^{\gamma}$ , thus
$\begin{aligned} Pr[\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y}]&=\sum_{\mathbf{z}\in\mathbb{Z}_{q}^{n}}Pr[\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y},\mathbf{b}=\mathbf{z}]\\ &=\sum_{\mathbf{z}\in S}Pr[\mathbf{x}=\mathbf{y}-\mathbf{A}\mathbf{z},\mathbf{b}=\mathbf{z}]\geq p\cdot 2^{\gamma}\end{aligned}.$
Thus it holds for all $\mathbf{z}\in\mathbb{Z}_{q}^{n}$ that
$Pr[\mathbf{b}=\mathbf{z}|\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y}]=\frac{Pr[\mathbf{b}=\mathbf{z},\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y}]}{Pr[\mathbf{A}\mathbf{b}+\mathbf{x}=\mathbf{y}]}\leq\frac{1}{2^{\gamma}}.$
B.3 Proof of L2FE-Hash Security
Proof of Theorem 1.
Let $D$ denote the procedure in Algo. 3 of obtaining a $d=(A,Ab+x)$ value from $x\stackrel{{\scriptstyle\mathcal{X}}}{{\leftarrow}}C_{\epsilon}$. From Theorem 11, we know that $\tilde{H}_{\infty}(X|D(X))\geq\gamma$. Note that $\tilde{H}_{\infty}(B|D(X))$ is the same value because for fixed $A,Ab+x$, there is a 1-to-1 correspondence between $b$ and $x$. Notice that $b\in\mathbb{Z}_{q}^{l}$ is a $l\log(q)$-bit string. Since Ext is an average case $(l\log(q),\gamma,\kappa,\epsilon_{\text{FE}})$-strong extractor, then by Definition 7, applying this strong extractor to $b$ gives $\textbf{SD}((\text{Ext}(B,K),D(X),K),(\mathcal{U}(\{0,1\}^{\kappa}),D(X),K))\allowbreak=\textbf{SD}((R,P),(\mathcal{U}(\{0,1\}^{\kappa}),P))\leq\epsilon_{\text{FE}}.$
Since Ext can extract $\kappa\leq\gamma-2\log(1/\epsilon_{\text{FE}})+2$ bits of randomness by Lemma 2.4 of [2004:fe], the proof completes. ∎
Appendix C FE Implements the Ideal Primitive
[…]
Definition 12 (HILL Entropy [1999:extractor-universal-hash-hill]).
[…]
Proof of Theorem 2.
Correctness proof. The correctness property of the ideal primitive is the same as that of FE.
Utility proof. The FE security property implies a utility argument that translates to the utility property Item 3 of the ideal primitive. Since $H_{\infty}(\mathcal{U}(\{0,1\}^{\kappa}),\mathcal{P})=H_{\infty}(\mathcal{U}(\{0,1\}^{\kappa}))=\kappa$, and $(\mathcal{R},\mathcal{P})\approx_{\epsilon_{\text{FE}}}(\mathcal{U}(\{0,1\}^{\kappa}),P)$ by FE security, then $H^{\text{HILL}}(\mathcal{R},P)\geq\kappa$ by Definition 12.
Security proof. To prove that FE security implies ideal primitive security, we first describe the FE and ideal primitive security games below. Subscripts on adversaries annotate the data utilized by them. The knowledge of input distribution $\mathcal{X}$ is assumed to be accessible to all adversaries.
Game $G_{0}$ (FE security game):
  • Challenger $\mathcal{C}$ flips a bit $b\leftarrow\mathcal{U}(\{0,1\})$.
    • If $b=1$, $\mathcal{C}$ samples $x\leftarrow\mathcal{X}$, and computes $(r,p)=\textsc{Gen}(x)$.
    • If $b=0$, $\mathcal{C}$ samples $x\leftarrow\mathcal{X}$, computes $(\_,p)=\textsc{Gen}(x)$, and samples $r\leftarrow\mathcal{U}(\{0,1\}^{\kappa})$.
  • Adversary $\mathcal{A}$ receives $(r,p)$ and outputs a guess $b^{\prime}$.
The game $G_{0}$ outputs $1$ if $b^{\prime}=b$. Advantage of $\mathcal{A}$ is defined as
$\text{Adv}_{0}(\mathcal{A})=\left|\Pr[\mathcal{A}_{(r,p,\mathcal{X})}=1\mid b=1]-\Pr[\mathcal{A}_{(p,\mathcal{X})}=1\mid b=0]\right|.$
Game $G_{1}$ (Ideal primitive security game):
  • Challenger $\mathcal{C}$ samples $x\leftarrow\mathcal{X}$ and computes $(r,p)=\textsc{Gen}(x)$.
  • Adversary $\mathcal{A}^{\prime}$ receives $(r,p)$ and outputs a guess $x^{\prime}$.
The game $G_{1}$ outputs $1$ if $d(x,x^{\prime})\leq t$. Advantage of $\mathcal{A}$ is
$\text{Adv}_{1}(\mathcal{A}^{\prime})=\left|\Pr[G_{1}(\mathcal{A}^{\prime}_{(r,p,\mathcal{X})})=1]-\Pr[G_{1}(\mathcal{A}^{\prime}_{(p,\mathcal{X})})=1]\right|.$
Game $G_{0}$ directly reduces to $G_{1}$ as shown in Algo. 4. In particular, the probability of $\mathcal{A}$ returning $1$ in either $b=0$ or $b=1$ case depends on the success probability of $\mathcal{A}^{\prime}$ in recovering $x^{\prime}$ and the correctness of Rep procedure. Therefore, we can derive $\text{Adv}_{0}(\mathcal{A})=(1-\delta)\text{Adv}_{1}(\mathcal{A}^{\prime})$:
1:Input: $(r,p)$
[…]
4:If $r^{\prime}=r$, then return $b^{\prime}:=1$, else return $b^{\prime}:=0$.
Algorithm 4 Construction of $\mathcal{A}$ with $\mathcal{A}^{\prime}$.
$\text{Adv}_{0}(\mathcal{A})$ $=\left|\Pr[\mathcal{A}_{(r,p,\mathcal{X})}=1\mid b=1]-\Pr[\mathcal{A}_{(p,\mathcal{X})}=1\mid b=0]\right|$
$=\left|\Pr[G_{1}(\mathcal{A}^{\prime}_{(r,p,\mathcal{X})})=1]-\Pr[G_{1}(\mathcal{A}^{\prime}_{(p,\mathcal{X})})=1]\right|$
$\,\cdot\Pr[\textsc{Rep}\text{ is correct}]$
$=(1-\delta)\text{Adv}_{1}(\mathcal{A}^{\prime})$
So $\epsilon_{\text{FE}}$}-FE security implies ${\epsilon_{\text{FE}}}(1-\delta)$-ideal primitive security ∎
Appendix D Prior Attacks Details
Attack details.
Bob attack uses an ID loss function which requires information about the feature extractor, hence making the attack whitebox [2024:bob]. Bob’s inversion model uses special blocks in its architecture called DSCasConv which are cascaded convolutions with skip connections.
GMI first trains a GAN model to learn the mapping between a latent space and the image data, then during the attack phase, uses the discriminator feedback and target model classification to optimize the latent vector corresponding to each user. KED-MI improves upon GMI by using the target model’s output labels in training of the GAN model, resulting in target-model specific attack models.
Implementation details.
For GMI and KED-MI attacks, we use the same attack pipeline as [2022:bido]. For GMI, we use their pretrained GAN model, and for KED-MI, we train our own GAN model for each target model. In particular, 1000 users were chosen to train the classification layer, with around 27 images per user in the train set, and 3 images per user in the test set. 30,000 images of 6453 users different from those 1000 chosen were used to train the KED-MI GAN model. For Bob, we use only one image per user to train the inversion model, for all 9177 users in our Celeb-A train set, without data augmentation.
To customize the attacks for target models with different protection schemes, we use the same model architecture as the facenet-pytorch model [2019:facenet-pytorch], and append a final layer that generates protection scheme outputs from the FaceNet embeddings. For example, the final layer returns $Rx||R.\text{flatten()}$ for the MRP protection scheme, $ss$ for Facial-FE, and $Ab+x||A.\text{flatten()}$ for L2FE-Hash.
Appendix E Supplementary Results
We evaluate the quality and similarity of reconstructed images to the original images by computing the Learned Perceptual Image Patch Similarity (LPIPS) metric [2018:LPIPS] and the Fréchet inception distance (FID) metric [2017:fid]. We claim that LPIPS is no different from Pipe ASR when using a different face recognition model during evaluation than the one used during training. Hence, we also evaluate Pipe ASR across 3 different feature extractor models. All images are resized to $64\times 64$ before evaluation.
LPIPS score.
LPIPS is a perceptual similarity metric developed as a proxy to human perception. It relies on the closeness of features obtained from a deep learning model. As seen in Table VIII, we observe two trends. First, we see that Pipe has a smaller score than Bob on average, indicating that Pipe reconstructed images are closer to the originals than Bob. Second, there is a sharp increase in the LPIPS score for L2FE-Hash. This is because while the reconstructed faces are close to human faces, they are not similar to the original face, indicating that L2FE-Hash is a successful defense.
FID score.
FID scores assess the quality of reconstructed images by comparing the distance between the reconstructed and original image distributions. We use the PyTorch-FID library implementation [2020:fid-implementation]. As seen in Table IX, FID score for Pipe is much lower than that of Bob, indicating that the distribution of images reconstructed using Pipe is of similar quality to that of original images. It is worth noting that the FID score increases significantly for Bob with MRP and L2FE-Hash protection schemes, unlike Pipe. This is because Bob is not able to produce good-quality face images given these embeddings, whereas PIPE still produces human-looking images that are not close to the original images. Hence, FID does not capture the privacy risk.
TABLE XIII: Cross model ASR (in %) of Pipe, trained and evaluated on the Celeb-A dataset. Column headings indicate the model used for evaluation.
Model Scheme Facenet ArcFace VGG-Face
Facenet Unprotected - 89.8 98.8
Facial-FE 87.8 97.8
MRP 76.1 88.2
L2FE-Hash 5.7 1.4
Arcface Unprotected 93.1 - 95.4
Facial-FE 90.2 91.3
MRP 36.7 41.8
L2FE-Hash 1.1 1.7
Cross-model Pipe ASR.
Since LPIPS relies on closeness of deep embeddings, we believe that Pipe ASR computed using a feature extractor different from the one used by the target model also represents a similarity metric. We use Facenet, ArcFace, and VGG-Face [2015:vgg-face] as feature extractors on the Celeb-A dataset to compute ASR.  Table XIII shows that the reconstructed images are still considered $t$-close to the original images and successfully authenticate with ASR $\geq 76\%$ in most cases. Therefore, the Pipe cross-model ASR is indicative of the performance we would expect from any other ML-based similarity metric like LPIPS.
Appendix F Reconstructed images
In this section we present some reconstructed images obtained by Bob and Pipe attack on the Celeb-A test dataset. We present reconstructed images for all setups– unprotected embeddings, FacialFE protected embeddings, MRP protected embeddings, and L2FE-Hash protected embeddings. We show the results for Facenet embeddings in Figure 4 and ArcFace embeddings in Figure 5. We observe that both Pipe and Bob can invert unprotected and FacialFE protected embeddings, though the images generated by Bob do not appear to be realistic. This is also reflected by the high FID score for Bob in Section 6. Unlike Bob, Pipe is also able to recover images somewhat similar to the original images from MRP protected embeddings. However, both Pipe and Bob on L2FE-Hash do not recover images similar to the original, demonstrating its security.
Figure 4: Images generated by attacking unprotected, FacialFE protected, MRP protected and L2FE-Hash protected Facenet embeddings. (a) Inversion using Pipe (b) Inversion using Bob. The first column in each subfigure shows the original image.
Figure 5: Images generated by attacking unprotected, FacialFE protected, MRP protected and L2FE-Hash protected ArcFace embeddings. (a) Inversion using Pipe (b) Inversion using Bob. The first column in each subfigure shows the original image.